Compute the uncollectible account expense, and make the appropriate journal entry

Johnson Corporation’s Unadjusted Trial Balance at year-end included the following accounts:

DebitCredit

Sales (75% represent creditsales)(credit)$1,152.000

AccountsReceivable(debit) $288,000

Allowance for DoubtfulAccounts(credit) $2,184


Compute the uncollectible account expense, and make the appropriate journal entry, for the current year assuming the uncollectible account expense is determined asfollows:

•a. Income statement approach, 1% of total sales.
•b. Income statement approach, 1.5% of credit sales.
•c. Balance sheet approach. The estimate based on an aging of accounts receivable is that an allowance of $12,000 would be appropriate.
Answer #1
Concepts and reason

Accounting: Accounting is a process of recording the transactions, classifying them in a specific manner, and is the process of summarizing and analyzing to interpret the results. It is a process of preserving the accounts.

Financial accounting is a process of preparing reports to provide all financial information to both the internal and external users of an organization. The financial statements that are prepared under the financial accounting are examined by independent certified public accountants, at the year-end, who express their opinion on the fairness of the reports shown by a company.

Income statement: This is the financial statement of a company that reports all the revenues that are earned and expenses that are to be expended by the company in the immediate accounting year. Income statement is also known profit and loss statement.

Balance sheet: Balance sheet refers to a statement of assets, liabilities, and owner’s equity as on a particular date of the fiscal year of the business enterprise. It also depicts the financial status of a business enterprise in a nutshell.

Transaction: Transaction is an act of buying or selling goods or rendering any service that is reliably measured in terms of money.

Journal entry: Journal entry is the recording of transactions in a systematic manner as they occur. Thus, it is a summary of all the transactions that has debit and credit aspects recorded chronologically.

Fundamentals

Allowance for doubtful debts: Allowance for doubtful debts is an account that is used to decrease the amount of accounts receivable. It indicates the estimation of company toward the account receivable.

Sales: Sales is an activity of selling the goods in the market that is sold by a seller and purchased by a buyer. It is the main source of revenue for the company. It is necessary to have consideration for sales.

Credit sales: Credit sales are the sales done, and the income or revenue is accrued but not received immediately. Cash is received during the future period for the sales. The customer is thus known as debtor.

Accounts payable: Accounts payable is the amount to be paid by a person or company who has purchased goods or received any services during the future period. It is the liability of the company and thus shown under liabilities in the balance sheet.

Accounts receivable: Accounts receivable is the amount to be received by a company that has sold goods or rendered any services during the period. It is the asset of the company and thus shown under assets in the balance sheet.

a)

Prepare the journal entry based on the income statement approach of total sales:

Credit
Debit
$11,520
Date Accounts title and explanation
Uncollectable expense
Allowance for doubtful accounts
(To record the

Therefore, the uncollectable expense is debited and the allowance for doubtful accounts is credited for $11,520.

Working notes:

Calculate the uncollectible expense:

Uncollectibleexpense=1%×Sales=1%×$1,152,000=$11,520\begin{array}{c}\\{\rm{Uncollectible expense = 1\% }} \times {\rm{Sales}}\\\\{\rm{ = 1\% }} \times \$ 1,152,000\\\\ = \$ 11,520\\\end{array}

Therefore, the uncollectible expense is $11,520.

b)

Prepare the journal entry based on the income statement approach of credit sales:

Date Accounts title and explanation
Uncollectable expense
Allowance for doubtful accounts
To record the uncollectable expense

Therefore, the uncollectable expense is debited and the allowance for doubtful accounts is credited for $12,960.

Working notes:

Calculate the uncollectible expense:

Uncollectableexpense=(Percentageofcreditsales×(Sales×Percentageofcreditsales))=1.5%×($1,152,000×75%)=1.5%×864,000=$12,960\begin{array}{c}\\{\rm{Uncollectable expense = }}\left( \begin{array}{l}\\{\rm{ Percentage of credit sales}}\\\\ \times \left( \begin{array}{l}\\{\rm{Sales}}\\\\ \times {\rm{Percentage of credit sales}}\\\end{array} \right)\\\end{array} \right)\\\\ = 1.5\% \times \left( {\$ 1,152,000 \times 75\% } \right)\\\\ = 1.5\% \times 864,000\\\\ = \$ 12,960\\\end{array}

Therefore, the uncollectible expense is $12,960.

c)

Prepare the journal entry based on the balance sheet approach:

Credit
Debit
$9,816
Date Accounts title and explanation
Uncollectable expense
Allowance for doubtful accounts
(To record the

Working notes:

Calculate the uncollectible expense:

Uncollectableexpense=(AccountsreceivablesAllowancefordoubtfulaccounts)=$12,000$2,184=$9,816\begin{array}{c}\\{\rm{Uncollectable expense = }}\left( \begin{array}{l}\\{\rm{Accounts receivables}}\\\\ - {\rm{Allowance for doubtful accounts}}\\\end{array} \right)\\\\ = \$ 12,000 - \$ 2,184\\\\ = \$ 9,816\\\end{array}

Therefore, the uncollectible expense is $9,816.

Ans: Part a

Date Accounts title and explanation
Uncollectable expense
Allowance for doubtful accounts
(To record the uncollectable expens
Answer #2

A. Income statement approach 1% of total sales.

Uncollectable Expense = 1%*1152000

Uncollectable Expense = 11,520

Journal Entry

Account Title Debit Credit
Uncollectable Expense 11520
Allowance for doubtful accounts 11520

B. Income statement approach 1.5% of credit sales.

Uncollectable Expense = 1.5%*(1152000*75%)

Uncollectable Expense = 12,960

Journal Entry

Account Title Debit Credit
Uncollectable Expense 12960
Allowance for doubtful accounts 12960

C. Balance sheet approach the estimate based on aging of accounts receivable is that an allowance of 12,000 would be appropriate.

Uncollectable Expense = 12000-2184

Uncollectable Expense = 9816

Journal Entry

Account Title Debit Credit
Uncollectable Expense 9816
Allowance for doubtful accounts 9816